# Modeling of AR 12760 with GX Simulator and Evidence for the Extended Transition Region in Peripheral Active Region Loops

### v1

@@ Abstract
In order to understand solar atmospheric heating it is important to test heating models against spatially resolved data from solar active regions. Here we model a small active region, AR 12760 observed on 2020 April 28, with the GX Simulator package by fitting the EUV intensities in wavebands observed by Solar Dynamics Observatory’s Atmospheric Imaging Assembly. We assume the temporally and spatially averaged heating rate along a loop has a power-law dependence on loop length, $L$ and average magnetic field strength along the loop, $B_{\mbox{\scriptsize avg}}$. We find that the best fit heating model for the 211 Å band is $\langle Q\rangle\approx 7\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{1.5}(L/{10^{9}\mbox{cm}})^{-1}$ erg cm-3 s-1 but that there is a range of parameters that give qualitatively reasonable fits, which we conclude is due to a correlation between $B_{\mbox{\scriptsize avg}}$ and $L$. In addition, we find that the models of the bands including cooler emission (131 and 171 Å) greatly underestimate the extent of the emission in the legs of the longer loops at the peripheries of the active region that are the strongest contributors of the emission in those bands. We conclude that this is because the modeling assumes that all transition region emission is confined to the loop foot points, but in reality the upper transition region of longer loops extends significantly farther into the loop. It is important to consider this aspect of the transition region in future efforts to model EUV emission.
@@ 1 Introduction
Active region modelers often take the approach of applying and constraining scaling laws describing heating to a set of model loops. There are many theoretically based relationships connecting the heating to quantities like loop geometry, magnetic field, and currents (Mandrini et al., 2000; Cranmer & Winebarger, 2019) to which the results can be compared (Gudiksen & Nordlund, 2002, 2005; Schrijver et al., 2004; Mok et al., 2005, 2008, 2016; Warren & Winebarger, 2006; Lundquist et al., 2008a, b; Ugarte-Urra et al., 2017) for the purpose of bolstering or discarding particular heating models. Of special interest to us is the use of scaling laws based explicitly on episodic “nanoflare” heating models, employed, for instance, by Barnes et al. (2019); Mondal et al. (2025).
@@
Another technique is to heat the corona using self-consistent MHD simulations in which coronal heating is produced by numerical resistivity and viscosity (Gudiksen & Nordlund, 2002, 2005; Bourdin et al., 2013; Rempel, 2017; Warnecke & Peter, 2019). This has the advantage of creating a self-consistent model, but the extent to which the heating in the simulation represents the actual heating on the Sun is not clear.
@@
In contrast to the works above that have their main (although not exclusive) focus on heating parameters, Plowman (2021, 2023) assumed fairly simple assumptions about the distribution of emission in loops so as to use the EUV emission to constrain the magnetic field.
@@
In this paper we focus on testing the uses of the GX Simulator model to simulate the emission of an active region observed in EUV in multiple wavebands by the Solar Dynamics Observatory’s Atmospheric Imaging Assembly (SDO/AIA) in order to determine the dependence of emissions on loop length, magnetic field strength and average heating rate. GX Simulator was originally developed for the purpose of modeling flare loops in microwaves and X-rays (Nita et al., 2015), but has since been expanded to model active regions in radio and EUV (Nita et al., 2018; Fleishman et al., 2021b; Nita et al., 2023; Fleishman et al., 2025). GX Simulator uses scaling-law based heating combined with magnetic field extrapolation calculated using a vector magnetogram (see Section 2). One advantage to GX Simulator over related approaches is that it computes the emission from every voxel in the volume modeled rather than along particular loops extrapolated upward from a grid of footpoints on the photosphere. Models of the latter type either have large emission-free gaps between loops in the corona or non-physical loop overlap.
@@
In Section 2 we describe the process used to model AR 12760 with GX Simulator, including the magnetic field, plasma properties, and fitting of the model to the EUV intensity distribution of the active region. Section 3 describes the fit results, with special focus on the 211 and 171 Å bands. In Section 4 we discuss the results including their indication of a relationship between the magnetic field and loop length, the importance of correctly modeling the upper transition region, and ways in which the modeling might be improved. In Section 5 we summarize our results.
2 Modeling Process
2.1 Overview
@@
GX Simulator is described extensively by Nita et al. (2018, 2023) and is available as the gx_simulator package in Solar Soft (Freeland & Handy, 1998). Here we discuss the points most relevant to this paper, starting with a very general overview and adding details in Sections 2.2-2.5.
@@ 2.2 EUV Data
We selected for study Active Region (AR) 12760, a small compact active region that crossed the meridian on 2020 April 28. It was chosen because of its compact structure, location near Sun center, lack of sunspots, and low level of activity. It was also the only significant active region on the Sun at the time, making it a good candidate for comparison with full-sun spectral irradiance measurements which, although not a part of this paper, is a longer term goal of a larger project.
@@
Figure 2: SDO (a)-(f) averages of AIA intensity for the 94, 131, 171, 193, 211, and 335 Å bands from 9:15-13:15 UT shown with linear scaling. (g) HMI intensity and (h) longitudinal magnetic field data at 11:11 UT.
@@
In order to constrain the heating parameters we used data from the 94, 131, 171, 193, 211, and 335 Å bands of Solar Dynamics Observatory’s Atmospheric Imaging Assembly (SDO/AIA; Pesnell et al. (2012); Lemen et al. (2012)). Because we wish to simulate the general, time averaged behavior of the active region rather than its emissions at any one moment, we used an average of the AIA data at a 5 minute cadence from 9:15-13:15 UT. This time period is slightly before the region crossed the meridian, but was selected to avoid small instances of emerging flux that occurred later in the day and could have affected the magnetic model. Each image was corrected for solar rotation before averaging, aligning it to the region’s location at 11:15 UT.
@@
At any given time, the observed active region corona is composed of a diffuse component and a component of bright loops (on the order of 1 Mm, or a few arc-seconds in width in the AIA images). There is more total emission in the diffuse component than the loop component, as loops are typically only modest (10-40%) enhancements over the background (Klimchuk & DeForest, 2020). These bright, thin loops appear and disappear at different locations with lifetimes of several thousand seconds. In a 4-hour average they tend to combine to form larger structures (with widths on the order of $10\arcsec$). A successful GX Simulator model would reproduce both the diffuse component and this medium-scale structure. Note that there is an abundance of theoretical evidence that both components are comprised of ultra-thin, spatially unresolved loops (e.g., Klimchuk et al., 2023; Johnston et al., 2025).
2.3 Magnetic Field Model
@@
To determine the coronal magnetic field we used the Automatic Model Production Pipeline (AMPP, Nita et al., 2023) initiated with the SDO/Helioseismic and Magnetic Imager (HMI; Scherrer et al. (2012)) vector magnetic field map from 11:12 UT.
@@
The magnetic field is calculated using a non-linear force free field (NLFFF) extrapolation as described in Fleishman et al. (2017) and based on the methods of Wiegelmann (2004). Following tests performed by Fleishman et al. (2017) we did not apply any preprocessing to the photospheric boundary condition. The C++ source code, along with set of compiled libraries for Windows, Linux, and Mac OS platforms, their calling IDL wrappers, and the GUI application, AMPP, are included in the GX Simulator SSW distribution package. These source codes and platform specific compiled libraries are automatically updated from their independently maintained GitHub development repository111 Magnetic-Field_Library. A Python utility based on the same library is also available from GitHub (pyAMPP, Yu et al., 2025). In addition, this NLFFF package may be directly downloaded from a Zenodo© digital repository (Stupishin, 2020) [^2] Magnetic Field Library: NLFFF and magnetic lines. The magnetic field strength in this active region is fairly low, making it somewhat difficult to model in detail, as will be discussed in Section 3.
@@
We developed two models, each of which is a compromise between resolution and field of view. Both models were centered at $x=-75\arcsec$, $y=-50\arcsec$ and consisted of cubes with 180 cells in the $x$ and $y$ directions and 200 cells in the $z$ direction. Magnetic Model 1 (Figure 3a) has a cell size of 1400 km so as to better model the core field lines but some of the longer loops extending from the active region periphery are “open” and thus not included in our modeling of the emission (see Section 2.4 below). Magnetic Model 2 (Figure 3b), conversely, covers an area twice as large while increasing the cell size to 2800 km to remain within memory constraints of the computer being used. Model 2 includes the long loops, but produces a more simplified model of the core of the active region. Sample field lines from both models are shown in Figure 3 in which field lines that close within the model box are shown in green, while ones that extend outside the box are in yellow.
2.4 Loop Plasma
The active region volume is divided into voxels and a value for $B_{\mbox{\scriptsize avg}}$ and $L$ is assigned to each voxel using the NLFFF model. Each voxel is characterized using the field line passing closest to its center, providing a value for $B_{\mbox{\scriptsize avg}}$ and $L$.
@@ 2.4 Loop Plasma
The plasma properties in the loop are determined using a look-up table as a function of $L$ and $\langle Q\rangle$, determined using the Enthalpy-Based Thermal Evolution of Loops (EBTEL, Klimchuk et al. (2008); Cargill et al. (2012a, b)) model with a specific distribution of heating events. EBTEL is a “0D” model that computes the evolution of the coronal-averaged temperature, pressure, and density in a loop. The instantaneous coronal DEM is assumed to be distributed uniformly over a temperature interval centered on the average temperature and having a width dictated by the expected ratio between the apex and average temperatures. In practice, the time-averaged coronal DEM is controlled primarily by loop evolution and not by the instantaneous distribution. The DEM is spread uniformly along the entire coronal portion of the loop. For the transition region EBTEL computes the detailed DEM distribution at each time in the simulation based on energy balance. However, all the DEM and therefore all the transition region emission is placed in a single cell at the loop footpoint.
@@
EBTEL takes the temperature boundary between the corona and transition region to be the value where thermal conduction switches from a cooling term above to a heating term below. This temperature is roughly 60% of the apex temperature. The lower and middle transition region by temperature are very thin geometrically, but 1D loop modeling indicates that the upper transition region extends an appreciable distance from the loop footpoint. This plays an important role in the model images, as we discuss later.
@@
Figure 4: Contours show regions in which it was assumed there was transition region emission using Magnetic Model 1 (top row) and Magnetic Model 2 (bottom row) superimposed on a HMI longitudinal magnetogram (left) and AIA 171 Å  image (right).
@@
In order to produce a lookup table of coronal and transition region DEMs as a function of $\langle Q\rangle$ and $L$, a large number of EBTEL runs are done. The grid provided logarithmically spaced values in the range $L/2=1$ to $6.3\times 10^{4}$ Mm. For each $L/2$ value there is a different range of $\langle Q\rangle$  with $\langle Q\rangle$ ranging from 0 to 794 erg cm-3 s-1 for $L/2=1$ and from $1.5\times 10^{-8}$ to 0.0126 for $L/2=6.3\times 10^{4}$ Mm. For each set of parameters a long run is averaged in order to represent a loop as a group of smaller strands heated randomly by an ensemble distribution of heating events. The currently available EBTEL lookup tables in GX Simulator are described in Nita et al. (2023). For this work we used ebtel_scale=1_alpha=-1.sav, which assumes a random distribution of nanoflares with an average time between nanoflares equal to the loop cooling time (scale = 1) and a power law distribution in the delay between events with index $\alpha=-1$. The intensity of the events is proportional to the time between them so that the intensity also has a power-law distribution with the same index. For each voxel the DEMs are determined from the EBTEL table using a bilinear interpolation. We assume a loop with uniform cross-sectional area. This can affect the emission and the ratio between coronal and transition region emission in particular (Cargill et al., 2022). The transition region is relatively fainter in expanding loops.
@@
Currently the way GX Simulator can model magnetic field lines that leave the model box is to either assume they contain plasma in hydrostatic equilibrium based on an assumed temperature and base pressure or to exclude them entirely. We have elected to leave out emission from these “open” field lines (open in the sense that they leave the box, not necessarily open to the solar wind). In addition, the emission was also set to zero for voxels centered on a small number of very long field lines not within the boundaries of the $\langle Q\rangle$,$L$ lookup table and in cases in which the fitting algorithm was attempting to use unrealistically low values of $q_{0}$ in an effort to find a good fit. The boundaries of the EBTEL table do not affect the results greatly in the parameter ranges with good fits to the data, although we discuss a possible issue related to open field lines in Sections 3.2 and 4.2.

### v2

@@ Abstract
In order to understand solar atmospheric heating it is important to test heating models against spatially resolved data from solar active regions. Here we model a small active region, AR 12760 observed on 2020 April 28, with the GX Simulator package by fitting the extreme-ultraviolet (EUV) intensities in wave bands observed by Solar Dynamics Observatory’s Atmospheric Imaging Assembly. We assume the temporally and spatially averaged heating rate along a loop has a power-law dependence on loop length, $L$ and average magnetic field strength along the loop, $B_{\mbox{\scriptsize avg}}$. We find that the best fit heating model for the 211 Å band is $\langle Q\rangle\approx 7\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{1.5}(L/{10^{9}\mbox{cm}})^{-1}$ erg cm-3 s-1 but that there is a range of parameters that give qualitatively reasonable fits, which we conclude is due to a correlation between $B_{\mbox{\scriptsize avg}}$ and $L$. In addition, we find that the models of the bands including cooler emission (131 and 171 Å) greatly underestimate the extent of the emission in the legs of the longer loops at the peripheries of the active region that are the strongest contributors of the emission in those bands. We conclude that this is because the modeling assumes that all transition-region emission is confined to the loop foot points, but in reality the upper transition region of longer loops extends significantly farther into the loop. It is important to consider this aspect of the transition region in future efforts to model EUV emission.
@@ I Introduction
Active region modelers often take the approach of applying and constraining scaling laws describing heating to a set of model loops. Here, “loop” refers to a magnetic flux tube, not necessarily a bright feature in an image. There is a diffuse component of the corona comprised of observationally indistinguishable loops. There are many theoretically based relationships connecting the heating to quantities like loop geometry, magnetic field, and currents (Mandrini et al., 2000; Cranmer and Winebarger, 2019) to which the results can be compared (Gudiksen and Nordlund, 2002, 2005; Schrijver et al., 2004; Mok et al., 2005, 2008, 2016; Warren and Winebarger, 2006; Lundquist et al., 2008a, b; Ugarte-Urra et al., 2017) for the purpose of bolstering or discarding particular heating models. Of special interest to us is the use of scaling laws based explicitly on episodic “nanoflare” heating models, employed, for instance, by Barnes et al. (2019); Mondal et al. (2025).
@@
Another technique is to heat the corona using self-consistent MHD simulations in which coronal heating is produced by numerical resistivity and viscosity (Gudiksen and Nordlund, 2002, 2005; Bourdin et al., 2013; Rempel, 2017; Warnecke and Peter, 2019). This has the advantage of creating a self-consistent model, sometimes called an ab initio model, which does not rely on assumptions about coronal heating. However, the drawback of this approach is that the resistivity and viscosity in the models are orders of magnitude larger than solar values, and the mechanism of heating may be fundamentally different from the true solar mechanism, e.g., relatively passive Ohmic dissipation versus explosive magnetic reconnection.
@@
In contrast to the works above that have their main (although not exclusive) focus on heating parameters, Plowman (2021, 2023) made fairly simple assumptions about the distribution of emission in loops so as to use the EUV emission to constrain the magnetic field.
@@
In this paper we focus on testing the uses of the GX Simulator model to simulate the emission of an active region observed in EUV in multiple wave bands by the Solar Dynamics Observatory’s Atmospheric Imaging Assembly (SDO/AIA) in order to determine the dependence of emissions on loop length, magnetic field strength and average heating rate. GX Simulator was originally developed for the purpose of modeling flare loops in microwaves and X-rays (Nita et al., 2015), but has since been expanded to model active regions in radio and EUV (Nita et al., 2018; Fleishman et al., 2021b; Nita et al., 2023; Fleishman et al., 2025a; Kuznetsov et al., 2025). GX Simulator uses scaling-law based heating combined with magnetic field extrapolation calculated using a vector magnetogram (see Section II). One advantage to GX Simulator over related approaches is that it computes the emission from every voxel in the volume modeled rather than along particular loops extrapolated upward from a grid of footpoints on the photosphere. Models of the latter type either have large emission-free gaps between loops in the corona or non-physical loop overlap at low altitudes.
@@
In Section II we describe the process used to model AR 12760 with GX Simulator, including the magnetic field, plasma properties, and fitting of the model to the EUV intensity distribution of the active region. Section III describes the fit results, with special focus on the 211 and 171 Å bands. In Section IV we discuss the results including their indication of a relationship between the magnetic field and loop length, the importance of correctly modeling the upper transition region, and ways in which the modeling might be improved. In Section V we summarize our results.
II Modeling Process
II.1 Overview
@@
GX Simulator is described extensively by Nita et al. (2018, 2023) and is available as the gx_simulator package in Solar Soft (Freeland and Handy, 1998). Here we discuss the points most relevant to this paper, starting with a very general overview and adding details in Sections II.2-II.5.
@@ II.2 EUV Data
We selected for study Active Region (AR) 12760, a small compact active region that crossed the meridian on 2020 April 28. It was chosen because of its compact structure, location near Sun center, lack of sunspots, and low level of activity. It was also the only significant active region on the Sun at the time, making it a good candidate for comparison with full-sun spectral irradiance measurements that, although not a part of this paper, pertains to a longer-term goal of a larger project.
@@
Figure 2: (a)-(f) averages of SDO/AIA intensity for the 94, 131, 171, 193, 211, and 335 Å bands from 9:15-13:15 UT shown with linear scaling. (g) HMI intensity and (h) longitudinal magnetic field data at 11:11 UT.
@@
In order to constrain the heating parameters we use data from the 94, 131, 171, 193, 211, and 335 Å bands of Solar Dynamics Observatory’s Atmospheric Imaging Assembly (SDO/AIA; Pesnell et al. (2012); Lemen et al. (2012)). Because we wish to simulate the general, time averaged behavior of the active region rather than its emissions at any one moment, we use an average of the AIA data at a 5 minute cadence from 9:15-13:15 UT (see Figure 2). This time period is slightly before the region crossed the meridian, but was selected to avoid small instances of emerging flux that occurred later in the day and could have affected the magnetic model. Each image was corrected for solar rotation before averaging, aligning it to the region’s location at 11:15 UT.
Figure 3: The 171 and 211 Å images at 11:15 UT (left) compared with the average images made using one image every five minutes from 09:15-13:15 UT (right). The on line-version of this paper includes an animation with the equivalent frame showing each image (one every 5 minutes from 09:15-13:15 UT) used to calculate the average. The video is 4 s long.
@@
At any given time, the observed active region corona is composed of a diffuse component and a component of bright loops (on the order of 1 Mm, or a few arcseconds in width in the AIA images). There is more total emission in the diffuse component than the loop component, as loops are typically only modest (10-40%) enhancements over the background (Klimchuk and DeForest, 2020). These bright, thin loops appear and disappear at different locations with lifetimes of several thousand seconds. Over a 4-hour average they tend to combine to form larger structures (with widths on the order of $10\arcsec$), as can be seen in Figure 3, which compares individual images taken at 11:15 UT with the average over 4 hours. A successful GX Simulator model would reproduce both the diffuse component and this medium-scale structure. Note that there is an abundance of theoretical evidence that both components are comprised of ultra-thin, spatially unresolved loops (e.g., Klimchuk et al., 2023; Johnston et al., 2025).
II.3 Magnetic Field Model
@@
To determine the coronal magnetic field we use the Automatic Model Production Pipeline (AMPP, Nita et al., 2023) initiated with the SDO/Helioseismic and Magnetic Imager (HMI; Scherrer et al. (2012)) vector magnetic field map from 11:12 UT.
@@
The magnetic field is calculated using a nonlinear force free field extrapolation as described in Fleishman et al. (2017) and based on the methods of Wiegelmann (2004). Following tests performed by Fleishman et al. (2017) we did not apply any preprocessing to the photospheric boundary condition. The top and side boundary conditions are taken from the initial potential-field extrapolation that is initiated by the $B_{z}$ bottom boundary condition. To minimize the effect of these boundary conditions on the NLFFF solution, we follow Wiegelmann (2004) and employ buffer zones at the top and side boundaries that comprise 10% of the corresponding size of the data cube as detailed in Nita et al. (2023). The C++ source code, along with set of compiled libraries for Windows, Linux, and Mac OS platforms, their calling IDL wrappers, and the GUI application, AMPP, are included in the GX Simulator SSW distribution package. These source codes and platform specific compiled libraries are automatically updated from their independently maintained GitHub development repository [^1] Magnetic-Field_Library. A Python utility based on the same library is also available from GitHub (pyAMPP, Yu et al., 2025). In addition, this NLFFF package may be directly downloaded from a Zenodo digital repository (Stupishin, 2023) [^2] Magnetic Field Library: NLFFF and magnetic lines. The magnetic field strength in this active region is fairly low, making it somewhat difficult to model in detail, as will be discussed in Section III.
@@
We developed two models, each of which is a compromise between resolution and field of view (FOV). Both models were centered at 2020 April 28 11:12 UT at $x=-75\arcsec$, $y=-50\arcsec$, and the point of view derives from those coordinates. The AIA data were adjusted to match that time. Thus, the view of the active region we are modeling is quite close to the actual view of AIA and not a simple top-down point of view. The models consisted of cubes with 180 cells in the $x$ and $y$ directions and 200 cells in the $z$ direction. Magnetic Model 1 (Figure 4a) has a cell size of 1400 km so as to better model the core field lines but some of the longer loops extending from the active region periphery are “open” and thus not included in our modeling of the emission (see Section II.4 below). Magnetic Model 2 (Figure 4b), conversely, covers an area twice as large while increasing the cell size to 2800 km to remain within the memory constraints of the computer being used. Model 2 includes the long loops, but produces a more simplified model of the core of the active region. Sample field lines from both models are shown in Figure 4 in which field lines that close within the model box are shown in green, while ones that extend outside the box are in yellow.
II.4 Loop Plasma
@@
The active region volume is divided into voxels and a value for $B_{\mbox{\scriptsize avg}}$ and $L$ is assigned to each voxel using the NLFFF model. Each voxel is characterized using the field line passing closest to its center, providing a value for $B_{\mbox{\scriptsize avg}}$ and $L$. The plasma properties in the loop are determined using a look-up table as a function of $L$ and $\langle Q\rangle$, determined using the Enthalpy-Based Thermal Evolution of Loops (EBTEL, Klimchuk et al. (2008); Cargill et al. (2012a, b)) model with a specific distribution of heating events. EBTEL is a “0D” model that computes the evolution of the coronal-averaged temperature, pressure, and density in a loop without spatial resolution. The instantaneous coronal DEM is assumed to be distributed uniformly over a temperature interval centered on the average temperature and having a width equal to twice the difference between the apex and average temperature. Because the time-averaged coronal DEM is determined by the behavior of many unresolved loops over time, it is reflective primarily of longer term loop evolution and not the instantaneous temperature distribution. Thus, we use the time average of long-duration EBTEL simulations that encompass many nanoflare energy releases. The time-averaged DEM is spread uniformly along the entire coronal portion of the loop. The EBTEL runs currently used to calculate GX Simulator input tables assume no loop expansion.
@@
For the transition region EBTEL computes a separate lookup table with the detailed DEM distribution at each time in the simulation based on energy balance and averages them to produce DEM curves for the $\langle Q\rangle$ and $L$ grid. EBTEL takes the temperature boundary between the corona and transition region to be the value where thermal conduction switches from a cooling term above to a heating term below. This temperature is roughly 60% of the apex temperature. The lower and middle transition region by temperature are confined to low altitudes, but 1D loop modeling indicates that the upper transition region extends an appreciable distance from the loop footpoint. This plays an important role in the model images, as we discuss in Section IV.2.
@@
Figure 5: Contours show regions in which it was assumed there was transition-region emission using Magnetic Model 1 (top row) and Magnetic Model 2 (bottom row) superimposed on a HMI longitudinal magnetogram (left) and AIA 171 Å image (right).
@@
In order to produce a lookup table of coronal and transition-region DEMs as a function of $\langle Q\rangle$ and $L$, a large number of EBTEL runs are done. The $60(Q)\times 49(L)$ grid provides logarithmically spaced values in the range $L/2=1$ to $6.3\times 10^{4}$ Mm. For each $L/2$ value there is a different range of $\langle Q\rangle$  with $\langle Q\rangle$ ranging from 0 to 794 erg cm-3 s-1 for $L/2=1$ and from $1.5\times 10^{-8}$ to 0.0126 for $L/2=6.3\times 10^{4}$ Mm. As discussed above, for each set of parameters a long run is averaged in order to represent a loop as a group of smaller strands heated randomly by an ensemble distribution of heating events. The currently available EBTEL lookup tables in GX Simulator are described in Nita et al. (2023). For this work we use ebtel_scale=1_alpha=-1.sav, which assumes a random distribution of nanoflares with an average time between nanoflares equal to the loop cooling time (scale = 1) and a power law distribution in the delay between events with index $\alpha=-1$. The energy of the events is proportional to the time between them so that the energy also has a power-law distribution with the same index.
GX Simulator calculates the transition region and corona emission independently and then sums them. For each voxel the DEMs are determined from the EBTEL table using a bilinear interpolation. GX Simulator places all the transition-region emission in a single layer at the bottom of the model cube. We assume a loop with uniform cross-sectional area. This can affect the emission and the ratio between coronal and transition-region emission in particular (Cargill et al., 2022). The transition region is relatively fainter in expanding loops, but how adding expansion would affect our model is a complex topic, as discussed in Section IV.3.
@@
Currently the way GX Simulator can model magnetic field lines that leave the model box is to either assume they contain plasma in hydrostatic equilibrium based on an assumed temperature and base pressure or to exclude them entirely. We have elected to leave out emission from these “open” field lines (open in the sense that they leave the box, not necessarily open to the solar wind). In addition, the emission was also set to zero for voxels centered on a small number of very long field lines not within the boundaries of the $\langle Q\rangle$,$L$ lookup table and in cases in which the fitting algorithm was attempting to use unrealistically low values of $q_{0}$ in an effort to find a good fit. The boundaries of the EBTEL table do not affect the results greatly in the parameter ranges with good fits to the data, although we discuss a possible issue related to open field lines in Sections III.2 and IV.2.